an e.m.f. represented by e=100 sin 100pi t is impressed across a circuit consisting of 40 ohms resistor in series with a 40microF capacitor and a 0.25 H indicator. Determine the r.m.s value of the current the power supplied the power factor.


1
Expert's answer
2021-04-19T17:10:06-0400

Resistance, "R" = "40" "\Omega"

Inductance, "L=0.25 \space H"

Capacitance, "C=40\space\mu F=40\times10^{-6}\space F"

"\omega=100\pi"

"X_L=\omega L" "=100\pi\times0.25=78.53"

"X_C=\dfrac{1}{\omega C}=\dfrac{1}{100\pi\times40\times10^{-6}}=79.57"

Impedance, "Z=\sqrt{R^2+(X_L-X_C)^2}=40.01\space\Omega"

Power factor, "cos\phi=\dfrac{R}{Z}=0.99"


"V_{RMS}=\dfrac{V_M}{2}=\dfrac{100}{2}=50\space V"

RMS value of current, "I_{RMS}=\dfrac{V_{RMS}}{Z}=\dfrac{50}{40.01}=1.24\space A"

Power Supplied, P"=I_{RMS}^2Zcos\phi=60.90\space W"